OpenStudy (anonymous):

A triangle has three vertices A(3,4), B(-3,2), C(-1,-4). Is this a right triangle. Explain. Please help been trying to solve this homework question for 20 min.!

OpenStudy (anonymous):

So, I don't know what to do from here or if it is right? m(AB) = (2-4)/ (-3-3) = -2/-6 = 1/3 m(BC) = (-4-2)/(-1+3)= -6/2 = -3 m(AC) = (-4-4)/ (-1-3) = -8/-4 = 2 m(AB)*m(AC) = 1/3 x 2 = 2/3

OpenStudy (anonymous):

Help!

OpenStudy (lucaz):

did you draw the trinagle with the points?

OpenStudy (lucaz):

I don't know how to explain this

OpenStudy (anonymous):

Is there anyone you know who does?

OpenStudy (lucaz):

people with high smartscore may know, but I don't know any to call =/

OpenStudy (wolf1728):

I can draw it

OpenStudy (candy13106):

ok well this one right

OpenStudy (candy13106):

ok well u have to first draw it

OpenStudy (candy13106):

r u there jackaroo

OpenStudy (anonymous):

yup i tried drawing still don't get it

OpenStudy (candy13106):

ok well wat grade r u in frist

OpenStudy (candy13106):

r u there

OpenStudy (anonymous):

no